after 12 years a man will be 3 times as old as he was 4 years ago. find his present age
Answer by stanbon(75887)

after 12 years a man will be 3 times as old as he was 4 years ago. find his present age
----
Let his current age be "x":
Equation:
x+12 = 3(x-4)
x + 12 = 3x -12
2x = 24
x = 12 yrs old
======================
Cheers,
Stan H.
======================